OpenStudy (anonymous):

What is the vertex of the graph of y = −3x^2 − x + 8?

OpenStudy (anonymous):

Step by step please? x :)

OpenStudy (anonymous):

step by step: the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) in your example \(a=-3,b=-1\)

OpenStudy (anonymous):

so what would you get for \(-\frac{b}{2a}\) if \(a=-3,b=-1\) ?

OpenStudy (anonymous):

-1/6

OpenStudy (anonymous):

ok good now for the second coordinate, replace \(x\) in \[y=-3x^2-x+8\] by \(-\frac{1}{6}\) and find \(y\) i.e. compute \[-3(-\frac{1}{6})^2-(-\frac{1}{6})+8\]

OpenStudy (anonymous):

the arithmetic is kind of annoying, but totally doable

OpenStudy (anonymous):

I think thats my problem, I'm not multiplying correctly :\

OpenStudy (anonymous):

\(-3\times \frac{1}{36}+\frac{1}{6}+8\) \(=-\frac{1}{12}+\frac{1}{6}+8\) \(=-\frac{1}{12}+\frac{2}{12}+8\) \(=\frac{1}{12}+8\)

OpenStudy (anonymous):

so either \(8\tfrac{1}{12}\) or \(\frac{97}{12}\)

OpenStudy (anonymous):

seem reasonable?

OpenStudy (anonymous):

The -1/12 part is whats kind of confusing me. It's probably dumb of me but how did you get it? I did everything else correctly

OpenStudy (anonymous):

ok so that part comes from \[-3\times \left(-\frac{1}{6}\right)^2\]

OpenStudy (anonymous):

square first, multiply second \[(-\frac{1}{6})\times( -\frac{1}{6})=\frac{1}{36}\]

OpenStudy (anonymous):

then \(-3\times \left(\frac{1}{36}\right)=\frac{-3}{36}\)

OpenStudy (anonymous):

reduce this fraction by dividing top and bottom by 3 and get \(-\frac{1}{12}\)

OpenStudy (anonymous):

hows that?

OpenStudy (anonymous):

That makes more sense, can we try another?

OpenStudy (anonymous):

sure i'm ready

OpenStudy (anonymous):

Where does the graph of y = 6x^2 − 11x − 10 cross the x-axis? I'm kinda rusty on it... -b/2a then plug in the numbers right?

OpenStudy (anonymous):

this is completely different because you are not asked for the vertex, you are asked for where it crosses the \(x\) axis it will cross the \(x\) axis if \(y=0\) so your job here is not to find the vertex, but rather to solve \[6x^2-11x-10=0\] for \(x\)

OpenStudy (anonymous):

fortunately this one factors as \[(2x-5)(3x+2)=0\]

OpenStudy (anonymous):

if you don't like factoring you can use the quadratic formula, but this one was cooked up to factor do you know how to solve \[(2x-5)(3x+2)=0\] for \(x\) ?

OpenStudy (anonymous):

Can you just keep going so I can see all of your steps, I get it but then again I don't so please continue :)

OpenStudy (anonymous):

ok so first you do understand i hope that you are not looking for the vertex, but rather you are looking to find the values of \(x\) for which \[6x^2-11x-10=0\] since this factors as \[(2x-5)(3x+2)=0\] we can set each factor equal to zero and solve

OpenStudy (anonymous):

if \(2x-5=0\) then \(2x=5\) and so \(x=\frac{5}{2}\) is one solution

OpenStudy (anonymous):

similarly solve \[3x+2=0\\3x=-2\\x=-\frac{2}{3}\]

OpenStudy (anonymous):

so it crosses the \(x\) axis at two point, \((-\frac{2}{3},0)\) and \((\frac{5}{2},0)\)

OpenStudy (anonymous):

this ok?

OpenStudy (anonymous):

Yeah, I'm going to study this. I'll be back in a few minutes, thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Part 1 − Find the vertex, axis of symmetry, domain, and range of the graph of y = −3x2 − 3x + 4. Show all work for full credit. Part 2 − Using complete sentences, explain how you can determine the axis of symmetry, the domain, and range without graphing y = −3x2 − 3x + 4. How do you find the domain & range

OpenStudy (anonymous):

did you repost this question?